Shift report 20011116 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Nov 16 2001 - 23:54:40 EST

  • Next message: Apache: "Shift report 20011117 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Nov 16 23:54:39 2001, Shift: 16:00-24:00
    Supervisor: rw, with: 
    Log book#10 pages 150 - 151
    
    Shift Summary:
    ------------------------------------------------------------
    1.Two storage at 16:20 and 21:30. ZDC rate in the beginning 
    was 500 and 1600, respectively. The first one was dumpoeted at 
    20:50. The second store has ZDC rate of 500 at the end 
    of the shift and still going.
    
    2.	run	trigger6 	Setup
    	5839	110K		FS ,MRS same sa the previous run
    	5840	100K            FS 1/8B d2=d3=d4=off MRS same
    	5841	 50K            FS  d1=d2=d3=d4=OFF  MRS same
    	5842	TPC pedestal
    	5843	DVM
    	5844 	Pedestals
    	5845	130K		FS 12deg-2/3A  MRS same(40-4A)
    	5846    120K            same sa above
    	5847    running 	same as above		
    
    3. No new problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Nov 16 2001 - 23:55:04 EST